Whispering Company purchased a new plant asset on April 1, 2020, at a cost of $744,000. It was estimated to have a service life of 20 years and a salvage value of $51,000. Whispering’ accounting period is the calendar year. Compute the depreciation for this asset for 2020 and 2021 using the sum-of-the-years'-digits method. (Round answers to 0 decimal places, e.g. 45,892.) Depreciation for 2020 $ enter a dollar amount Depreciation for 2021 $ enter a dollar amount eTextbook and Media Compute the depreciation for this asset for 2020 and 2021 using the double-declining-balance method. (Round answers to 0 decimal places, e.g. 45,892.) Depreciation for 2020 $ enter a dollar amount Depreciation for 2021 $ enter a dollar amount
Answer #1

Sum of years digits method,

>> Depreciable asset = $ 744,000 - $ 51,000 = $ 693,000.

>> Sum of years = n ( n + 1 ) / 2 = 20 ( 20 + 1 ) / 2 = 210.

>> Depreciation for 2020 = $ 693,000 * ( 20 / 210 ) * ( 9 / 12 )

>> Depreciation for 2020 = $ 49,500..

>> Depreciation for 2021 = $ 693000 * ( 19.25 / 210 )

>> Depreciation for 2021= $ 63,525.

======================================================================================

Double declining method.

>> Depreciation rate = 200 / Useful life = 200 / 20.

>> Depreciation rate = 10 %.

>> Depreciation for 2020 = $ 693,000 * 10 % * ( 9 / 12 )

>> Depreciation for 2020 = $ 51,975.

>> Depreciation for 2021 = ( $ 693,000 - $ 51,975 ) * 10 %.

>> Depreciation for 2021 = $ 64,103

========================================================================================